OpenStudy (anonymous):

SOLVING EXPONENTIAL EQUATIONS WITH UNEQUAL BASES: 5^(x+5)=9^x

OpenStudy (anonymous):

x ≈ –3.63 x ≈ –0.28 x ≈ 0.07 x ≈ 13.69

OpenStudy (anonymous):

\(\large 5^{x+5}=9^x \) \(\large 5^x\cdot5^5=9^x \) \(\large 5^5=\frac{9^x}{5^x}=(\frac{9}{5})^x \) \(\large 5^5=(\frac{9}{5})^x \rightarrow x=\frac{ln5^5}{ln(\frac{9}{5})} \)

OpenStudy (anonymous):

use your calculator from here...